OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 6 - x = 5x + 30 4x - 8 = 3x + 2 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (texaschic101):

6 - x = 5x + 30 -- subtract 5x from both sides 6 - x - 5x = 30 -- subtract 6 from both sides -x - 5x = 30 - 6 -- combine like terms -6x = 24 x = 24/-6 x = - 4 4x - 8 = 3x + 2 -- subtract 3x from both sides 4x - 3x - 8 = 2 -- add 8 to both sides 4x - 3x = 8 + 2 -- combine like terms x = 10 Eq 1 is smaller then eq 2